OpenStudy (andpeggy):

(I just need help setting this up please!) A confectioner is going to mix candy worth $3.90 per pound with candy worth $2.50 per pound to obtain 70 pounds of candy worth $3.30 per pound. How many pounds of each kind should she use?

satellite73 (satellite73):

want to do it with numbers?

OpenStudy (misty1212):

ok lets do it with numbers...

satellite73 (satellite73):

say you use 20 pounds of the $3.90 candy. how many pounds of the $2.50 would you use?

OpenStudy (misty1212):

idk

satellite73 (satellite73):

hint, the total is 70 pounds, you used 20

OpenStudy (misty1212):

oh ok 50

satellite73 (satellite73):

right , and what would the cost of the $3.90 candy be total?

OpenStudy (misty1212):

um \(3.90\times 20\)?

satellite73 (satellite73):

yea right aka \(\$78\)

OpenStudy (misty1212):

oh ok and the 50 pounds at $2.50 would cost $125 right?

satellite73 (satellite73):

yeah, \(2.50\times 50=125\) so what would the total be?

OpenStudy (misty1212):

add right? 78+125=203

satellite73 (satellite73):

yes, and 203 at 70 pounds would be what ?

OpenStudy (misty1212):

203/70=2.9?

OpenStudy (misty1212):

so what do we do now?

satellite73 (satellite73):

try again, this time with variables say you use \(x\) amount of \(\$3.90\) candy, how much of the \(\$3.90\) candy do you use?

OpenStudy (misty1212):

??

satellite73 (satellite73):

how did you get the 50?

OpenStudy (misty1212):

oh subtract \(70-x\) right?

satellite73 (satellite73):

yes subtract now the \(x\) pounds of \(\$3.90\) candy cost \(3.90\times x\) how much does the other candy cost?

OpenStudy (misty1212):

must be \(2.50\times (70-x)\) ?

satellite73 (satellite73):

right, and the total has to be \(70\times 3.30\) so set \[3.9x+2.5\times (70-x)=70\times 3.3\] and solve for \(x\)

OpenStudy (misty1212):

ooh ok

OpenStudy (misty1212):

i get it thanks !

OpenStudy (misty1212):

cant you do this some other way?

satellite73 (satellite73):

sure, the math teacher way say \[x+y=70\\ 3.9x+2.5y=3.3\times 70\] but all the work is teh same

OpenStudy (mathnerdz):

@satellite73 you are outstanding!!! :)